3/4X-1/4X+7-2=4
10(b-3)=2(6+b)

Answer by farriem(7)

1) Simplify
2) Subtract 5 from both sides
3) Mutiply both sides by 2
4) Check
Work
1) 3/4x-1/4x+7-2=4
2/4x+5=4
1/2x+5=4
2) 1/2x+5-5=4-5
1/2x=-1
3) 1/2x*2=-1*2
x=-2
4) 3/4(-2)-1/4(-2)+7-2=4
-6/4- -2/4 +7-2=4
-1 1/2 +1/2 +7-2=4
-1+7-2=4
4=4